OpenStudy (korosh23):

**Physics 12** 3. A student has a weight of 655 N. While riding on a roller-coaster this same student has an apparent weight of 1.96 x 103 N at the bottom of the dip that has a radius of 18.0 m. What is the speed of the roller-coaster? Answer: (18.76 m/s)

OpenStudy (korosh23):

You have the answer. Please, show me how to get that answer. Thank you

OpenStudy (vincent-lyon.fr):

Well, if you are requesting help, you should post your own attempt here first.

OpenStudy (michele_laino):

hint: the mass \(m_0\) of the student, is: \[\Large m_0 = \frac{P}{g} = \frac{{655}}{{9.81}} = ...?\] now, the apparent acceleration \(a_C\), is: \[\Large {a_C} = \frac{{{P_{apparent}}}}{m} = \frac{{1.96 \cdot {{10}^3}}}{{{m_0}}} = ...?\] and the magnitude of such apparent acceleration, has to be equal to the centripetal acceleration: \[\Large {a_C} = {a_{centripetal}} = \frac{{{v^2}}}{{18}} = ...?\] where \(v\) is the requested speed

OpenStudy (korosh23):

Yes, now I understand it. Apparent weight means force of normal, and I did not know it.

OpenStudy (korosh23):

Since I need the free body diagram first.

OpenStudy (korosh23):

Thank you

OpenStudy (vincent-lyon.fr):

Exactly, apparent weight is what you would read on a weighing scale giving out newtons.